# Electron-phonon coupling in magnetic materials using the local spin density approximation

### v1

@@ Abstract
Magnetic materials are crucial for manipulating electron spin and magnetic fields, enabling applications in data storage, spintronics, charge transport, and energy conversion, while also providing insight into fundamental quantum phenomena. In numerous applications, the interaction between electrons and lattice vibrations, known as electron-phonon coupling, can be of significant importance. In that regard, we extend the EPW package to be able to interpolate the electron-phonon matrix elements combining perturbation theory and maximally localized Wannier functions. This allows to use dense momentum grids at a reasonable computational cost when computing electron-phonon-related quantities and physical properties. We validate our implementation considering ferromagnetic iron and nickel, where we explore the absence of phonon-driven superconductivity, finding that superconductivity is intrinsically suppressed. Furthermore, we evaluate the carrier resistivity at finite temperatures for both systems, considering the role of the magnetic phase in carrier transport. Our findings indicate that in the case of Fe, the primary contributor to resistivity is electron-phonon scattering. In contrast, for Ni, electron-phonon scattering constitutes less than one-third of the resistivity, underscoring a fundamental difference in the transport properties of the two systems.
@@ I Introduction
Magnetic materials play an important role for technological applications, including magnetic cores in electric motors and generators, or magnetic memory devices in computers and mobile phones. From a historical standpoint, iron is arguably the most significant magnetic material. Iron can be alloyed to produce metals, insulators [1], superconductors [2], strong ferromagnetism [3], antiferromagnetism [4], thermal shielding at low and high temperatures and has good mechanical properties [5].
@@
Interest in magnetic materials is at an all-time high due the end of Moore’s law [6] and other limitations that current electronic devices are facing. In that regard, magnetic materials are seen as an appealing solution leveraging the spin of the electron giving rise to the field of spintronics. Spintronic devices exploit both the electronic and spin degrees of freedom [7, 8, 9], offering promising logic and memory architectures with inherently low power consumption [10, 11, 12, 13]. In addition, spintronic devices are widely used for memory applications [14, 15, 16, 17]. Iron and nickel are ubiquitous in steels, alloys, and functional magnetic devices used in the energy sector. Thus, there is renewed effort to understand, search and characterize potential materials relevant for the mentioned applications [18, 19, 20, 21]. Apart from these advances in devices for classical computers, the field of quantum computing might also require magnetic compounds that present specific transport properties [22, 23], and certain robustness to thermal noise to build and shield the quantum bits [24, 25].
@@
Given that the electrical and thermal conductivity in metals and semiconductors is strongly influenced by electron–phonon scattering, developing accurate first-principles methods to capture this mechanism is central to understanding and reducing energy losses in conductive materials. Although electron-phonon interactions have been studied for decades, accurately describing them has proven to be a challenging task. The situation becomes even more complicated when one considers the effect that phonons can have on the local magnetic structure [26, 27, 28, 29, 30], where in some cases they can even induce a non-zero transient magnetization in nonmagnetic materials [31, 32]. As a result, it is common to rely on effective models that take input from experiments [33, 34, 35], or previously computed ab initio quantities [36, 37, 38, 39]. Today, first-principles electron-phonon interactions are widely available thanks to density functional theory (DFT) [40, 41] and density functional perturbation theory (DFPT) [42, 43, 44, 45]. However, this approach is not free from challenges that can make actual calculations impractical. One of the main bottlenecks is the high computational cost required to sample the Brillouin zone in both electronic $\mathbf{k}$ and phononic $\mathbf{q}$ momentum space when integrating the different observables. In that regard, one of the most widespread solutions is to combine DFPT with maximally localized Wannier functions [46, 47] which enables interpolation at arbitrary momentum $\mathbf{k}$ and $\mathbf{q}$ of the electron-phonon matrix elements (EPME) $g_{mn\nu}(\mathbf{k},\mathbf{q})$, electronic eigenvalues $\varepsilon_{n\mathbf{k}}$ and eigenvectors $\psi_{n\mathbf{k}}$. Most implementations to date assume time-reversal symmetry. This assumption is only suited to study non-magnetic materials, hindering a global understanding of electron-phonon interactions. In fact, this is an important barrier to understanding the role of phonons in spintronics, caloritronics, or superconductivity and charge density waves in the presence of magnetism. More importantly, both electron and phonon dispersions can be affected by including the spin degree of freedom in the calculations while retaining the same exchange and correlation functional [48]. For example, neglecting the spin of iron leads to unstable phonons [49].
@@
This limitation has been addressed using finite differences [48, 50, 51, 52, 53, 54, 55] but is limited due to the computational cost of supercells. Recently, support for DFPT with non-collinear magnetism and ultrasoft pseudopotentials has been implemented in Abinit [56, 57, 58, 49, 59] and Quantum ESPRESSO [60, 61, 62]. However, it has never been extended to the interpolation of EPMEs, which hinders realistic applications in magnetic materials. We present a first-principles implementation for electron–phonon properties in magnetic systems, validated and applied to Fe and Ni. Our results reveal the essential role of the magnetic phase in resistivity, the microscopic origin of scattering in different ferromagnets, and the absence of phonon-mediated superconductivity in both materials. These insights provide a foundation for designing magnetic materials with tailored transport properties for energy applications. In this work, we extend the capabilities of the EPW code [46, 63, 64] to support collinear magnetism, unlocking the path to study electron-phonon related properties in magnetic materials with high accuracy and at reasonable computational cost. To that end, we exploit the recent developments in the Abinit code on spin-dependent electron-phonon interactions [57] to validate our implementation. This choice is guided by the documented small discrepancy between Abinit and Quantum ESPRESSO [65, 66, 67]. Although collinear magnetism serves as an approximation to its non-collinear counterpart when considering spin-orbit coupling (SOC), its computational efficiency is markedly superior, while retaining a good degree of accuracy.
@@
We study the carrier transport properties of ferromagnetic (FM) body-centered cubic (BCC) Fe and face-centered cubic (FCC) Ni when electron phonon interactions are accounted. Our findings indicate that the omission of magnetism in Fe leads to the emergence of imaginary (soft) phonon modes, confirming previous works. As sometimes done in the literature, excluding imaginary phonon modes overestimated resistivity due to the large contribution of modes on the onset of being imaginary. In contrast, Ni is a weaker ferromagnet with similar ferromagnetic and non-magnetic phonon band dispersions but a very different change of resistivity with temperature. Overall, the inclusion of magnetism reduces resistivity by a factor of three and seven for Fe and Ni, respectively. Finally, we compute the phonon-driven superconducting transition temperature ($T_{\rm c}$) in Fe and Ni and find vanishingly small values, which shows that neither material would be superconducting, even if their Cooper pairs could survive in the FM magnetization field. This work opens the gate towards accounting for the role of phonons in magnetic systems, and possible superconducting states in magnetic materials (such as iron arsenides and possibly cuprates), which are highly relevant for new energy-efficient technologies and energy loss in conducting materials. This work is organized as follows: we first present the theory in Sect. II, details about our computational framework in Sect. III, the interpolation of basic electron, phonon and electron-phonon quantities in Sect. IV, the calculation of physical properties and electron-phonon properties in Sect. V, and the final remarks and conclusions in Sect. VI.
@@ II Theory
where $\sigma$ and $\sigma^{\prime}$ are the spinor indices. From the definition of $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$, we can identify two terms. The first is proportional to $\omega_{\mathbf{q}\nu}^{-\frac{1}{2}}$. The second is the spin-dependent deformation potential $D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ which can be computed as [69]
@@
with $m_{\mathrm{u}}$ being the atomic mass constant. $D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ gives a measure of how much a phonon perturbation has changed the internal self-consistent potential felt by the electrons. In taking the absolute value of $|g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})|$, we take the root mean square average over any degenerate manifolds in $m$, $n$ and $\nu$. Note that the definitions of $D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ and $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ are also valid for the case of nonmagnetic compounds with and without SOC. The main property exploited in the nonmagnetic case are the following extra symmetries in the spin indices $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})]^{*}=(-1)^{\sigma+\sigma^{\prime}}g^{\bar{\sigma}^{\prime}\bar{\sigma}}_{mn\nu}(\mathbf{-k},\mathbf{-q})$, with $\bar{\sigma}$ being the opposite spin to $\sigma$. This means that if $\sigma=\sigma^{\prime}$ then $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})]^{*}=g^{\bar{\sigma}^{\prime}\bar{\sigma}}_{mn\nu}(\mathbf{-k},\mathbf{-q})$, and if $\sigma\neq\sigma^{\prime}$ then $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})]^{*}=-g^{\bar{\sigma}^{\prime}\bar{\sigma}}_{mn\nu}(\mathbf{-k},\mathbf{-q})$ (e.g. for heavy elements with spin orbit coupling but no magnetization). These two results are a consequence of time-reversal symmetry, which is absent in magnetic compounds.
In practice, performing magnetic calculations with SOC, commonly known as non-collinear magnetic calculations, is computationally expensive (typically by a factor of $2^{3}=8$ for cubic scaling DFT) in systems with several atoms. The collinear approximation consists in separating the two spin channels into up ($\uparrow$) and down ($\downarrow$) without explicit interaction between them, thereby only doubling the cost of the calculations, regardless of the system size. This allows us to define a spin-resolved EPME:
@@
In systems with many valence electrons, we are commonly interested in the deformation potential of several bands. In that regard, for each spin channel we can consider
@@
as the deformation potential of interest, with the sum over the electronic bands $m,n$ restricted to the bands in the Wannier manifold $\mathcal{M}$. We also extend the mode- and momentum-resolved electron-phonon coupling $\lambda_{\mathbf{q}\nu}$, which in the double delta approximation at the collinear level becomes [70]
$\lambda_{\mathbf{q}\nu}\equiv\sum_{\sigma}\lambda_{\mathbf{q}\nu}^{\sigma}=\sum_{\sigma mn\mathbf{k}}\frac{w_{\mathbf{k}}}{N(\varepsilon_{\rm F})\omega_{\mathbf{q}\nu}}|g_{mn\nu}^{\sigma}(\mathbf{k},\mathbf{q})|^{2}\\ \times\delta(\varepsilon_{\rm F}-\varepsilon^{\sigma}_{n\mathbf{k}})\delta(\varepsilon_{\rm F}-\varepsilon^{\sigma}_{m\mathbf{k+q}}),$ (7)
where $N(\varepsilon_{\textrm{F}})$ is the total density of states at the Fermi energy $\varepsilon_{\rm F}$ and $w_{\mathbf{k}}$ are the $\mathbf{k}$-point weights normalized to one. We note that another definition exists for the spin-resolved electron-phonon coupling using the partial density of states at the Fermi level $N^{\sigma}(\varepsilon_{\rm F})$ [58]. However, this convention will suffer from divergences in half- and semi-metals, where only one spin channel contributes to the total density of states. The present version should be used systematically for spin polarized cases, heeding the factor of 2 between $N^{\sigma}$ and $N$ in the unpolarized case.
Another important quantity is the mode and momentum resolved phonon linewidth $\gamma_{\mathbf{q}\nu}$
@@
where $w_{\mathbf{q}}$ are the $\mathbf{q}$-point weights normalized to one. The Eliashberg spectral function is used to determine the contribution of specific phonon modes to the total electron-phonon coupling strength defined as
@@
and can be used in the context of superconductivity to evaluate the critical temperature $T_{\rm c}$ assuming a phonon-mediated pairing mechanism and the Allen-Dynes formula [71]:
$T_{\rm c}^{\rm AD}=\frac{\hbar\omega_{\rm log}}{1.2k_{\rm B}}\mathrm{exp}{\left(\frac{-1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\right)},$ (11)
where $k_{\rm B}$ is the Boltzmann constant, $\mu^{*}$ is the semi-empirical Coulomb potential, and $\omega_{\rm log}$ is the logarithmic average of the phonon frequency (geometric average) defined as
[…]
Although this formula is generally used for $s$-wave pairing mechanism with time reversal symmetry, a BCS type of formula can be obtained for all the other gap channels $p$, $d$, or $f$ with spin singlet and triplet type of pairing (breaking time reversal) [72, 73]. Although certain deviations are expected, one can still explore the possibility of this pairing mechanism in magnetic compounds by means of the Allen-Dynes formula.
@@
where $f_{m\mathbf{k+q}}^{\sigma}$ is the Fermi-Dirac occupation for the spin $\sigma$ and $n_{\mathbf{q\nu}}$ is the Bose-Einstein distribution function. As for previous quantities, the average scattering rate is additive in the collinear approximation to magnetism:
@@
Prior studies have used the lowest-order variational approximation (LOVA) to study the resistivity of magnetic systems [49] which is based on computing the average scattering rate. This is a strong approximation in anisotropic systems or systems with different Fermi sheets [74]. Here, we solve the linearized Boltzmann transport equation (BTE) to calculate the conductivity [75, 76]
@@
which is the usual law of addition for resistivities in parallel, and in this case $\rho_{\alpha\beta}^{\sigma}$ is the resistivity experienced by the electrons in each spin channel. When comparing to polycrystalline samples, we use
@@
$\partial_{E_{\beta}}f_{n\mathbf{k}}^{\sigma}=ev_{n\mathbf{k}\beta}^{\sigma}\frac{\partial f^{\sigma}_{n\mathbf{k}}}{\partial\varepsilon^{\sigma}_{n\mathbf{k}}}\tau_{n\mathbf{k}}^{\sigma}+\frac{2\pi\tau_{n\mathbf{k}}^{\sigma}}{\hbar}\sum_{\mathbf{q}m\nu}w_{\mathbf{q}}|g_{mn\nu}^{\sigma}(\mathbf{k},\mathbf{q})|^{2}\\ \times[(n_{\mathbf{q\nu}}+1-f_{m\mathbf{k+q}}^{\sigma})\delta(\varepsilon_{n\mathbf{k}}^{\sigma}-\varepsilon_{m\mathbf{k+q}}^{\sigma}-\hbar\omega_{\mathbf{q}\nu})\\ \!+\!(n_{\mathbf{q\nu}}\!+\!f_{m\mathbf{k+q}}^{\sigma})\delta(\varepsilon_{n\mathbf{k}}^{\sigma}\!-\!\varepsilon_{m\mathbf{k+q}}^{\sigma}\!+\!\hbar\omega_{\mathbf{q}\nu})]\partial_{E_{\beta}}f_{m\mathbf{k+q}}^{\sigma}.$ (18)

### v2

@@ Abstract
Magnetic materials are crucial for manipulating electron spin and magnetic fields, enabling applications in data storage, spintronics, charge transport, and energy conversion, while also providing insight into fundamental quantum phenomena. In numerous applications, the interaction between electrons and lattice vibrations, known as electron-phonon coupling, can be of significant importance. In that regard, we extend the EPW package to be able to interpolate the electron-phonon matrix elements combining perturbation theory and maximally localized Wannier functions. This enables the use of dense momentum grids at a reasonable computational cost when computing electron-phonon-related quantities and physical properties. We validate our implementation considering ferromagnetic iron and nickel, where we explore the absence of phonon-driven superconductivity, finding that superconductivity is intrinsically suppressed. Furthermore, we evaluate the carrier resistivity at finite temperatures for both systems, considering the role of the magnetic phase in carrier transport. Our findings indicate that in the case of Fe, the primary contributor to resistivity is electron-phonon scattering. In contrast, for Ni, electron-phonon scattering constitutes less than one-third of the resistivity, underscoring a fundamental difference in the transport properties of the two systems.
@@ I Introduction
Magnetic materials play an important role for technological applications, including magnetic cores in electric motors and generators, or magnetic memory devices in computers and mobile phones. From a historical standpoint, iron is arguably the most significant magnetic material. Iron can be alloyed or combined with other elements to form compounds that are metals, insulators [4], superconductors [58], strong ferromagnetism [66], antiferromagnetism [79], thermal shielding at low and high temperatures and has good mechanical properties [59].
@@
Interest in magnetic materials is at an all-time high due to the end of Moore’s law [104] and other limitations that current electronic devices are facing. In that regard, magnetic materials are seen as an appealing solution leveraging the spin of the electron giving rise to the field of spintronics. Spintronic devices exploit both the electronic and spin degrees of freedom [124, 54, 17], offering promising logic and memory architectures with inherently low power consumption [34, 111, 11, 32]. In addition, spintronic devices are widely used for memory applications [56, 2, 13, 12]. Iron and nickel are ubiquitous in steels, alloys, and functional magnetic devices used in the energy sector. Thus, there is renewed effort to understand, search and characterize potential materials relevant for the mentioned applications [6, 61, 77, 39]. Apart from these advances in devices for classical computers, the field of quantum computing might also require magnetic compounds that present specific transport properties [26, 55], and a certain robustness to thermal noise to build and shield the quantum bits [31, 97].
@@
Given that electrical and thermal conductivity in metals and semiconductors is strongly influenced by electron–phonon scattering, developing accurate first-principles methods to capture this mechanism is central to understanding and reducing energy losses in conductive materials. Although electron-phonon interactions have been studied for decades, accurately describing them has proven to be a challenging task. The situation becomes even more complicated when one considers the effect that phonons can have on the local magnetic structure [102, 84, 105, 1, 33], where in some cases they can even induce a non-zero transient magnetization in nonmagnetic materials [51, 10]. As a result, it is common to rely on effective models that take input from experiments [99, 101, 19], or previously computed ab initio quantities [35, 37, 62, 63]. Today, first-principles electron-phonon interactions are widely available thanks to density functional theory (DFT) [57, 65] and density functional perturbation theory (DFPT) [9, 47, 98, 8]. However, this approach is not free from challenges that can make actual calculations impractical. One of the main bottlenecks is the high computational cost required to sample the Brillouin zone in both electronic $\mathbf{k}$ and phononic $\mathbf{q}$ momentum space when integrating the different observables. In this regard, one of the most widespread solutions is to combine DFPT with maximally localized Wannier functions [42, 76], which enables interpolation at arbitrary momentum $\mathbf{k}$ and $\mathbf{q}$ of the electron-phonon matrix elements (EPME) $g_{mn\nu}(\mathbf{k},\mathbf{q})$, electronic eigenvalues $\varepsilon_{n\mathbf{k}}$ and eigenvectors $\psi_{n\mathbf{k}}$. Most implementations to date assume time-reversal symmetry. This assumption is only suited to the study of non-magnetic materials, hindering a global understanding of electron-phonon interactions. In fact, this is an important barrier to understanding the role of phonons in spintronics, caloritronics, or superconductivity and charge density waves in the presence of magnetism. More importantly, both electron and phonon dispersions can be affected by including the spin degree of freedom in the calculations while retaining the same exchange and correlation functional [21]. For example, neglecting the spin of iron leads to unstable phonons [73].
@@
This limitation has been addressed using finite differences [25, 36, 21, 22, 23, 114, 121, 89, 115] but is limited due to the computational cost of supercells. Recently, support for DFPT with non-collinear magnetism and ultrasoft pseudopotentials has been implemented in Abinit [45, 46, 113, 73, 112] and Quantum ESPRESSO [41, 40, 106], as well as interpolation of the electron-phonon matrix elements with collinear magnetism in Perturbo [122, 123, 24, 68]. We here present a first-principles implementation for electron–phonon properties in magnetic systems, validated and applied to Fe and Ni. Our results reveal the essential role of the magnetic phase in resistivity, the microscopic origin of scattering in different ferromagnets, and the absence of phonon-mediated superconductivity in both materials. These insights provide a foundation for designing magnetic materials with tailored transport properties for energy applications. In this work, we extend the capabilities of the EPW code [42, 91, 69] to support collinear magnetism, unlocking the path to study electron-phonon related properties in magnetic materials with high accuracy and at reasonable computational cost. To this end, we exploit recent developments in the Abinit code on spin-dependent electron-phonon interactions [46] to validate our implementation. This choice is guided by the documented small discrepancy between Abinit and Quantum ESPRESSO [90, 16, 93]. Although collinear magnetism serves as an approximation to its non-collinear counterpart when considering spin-orbit coupling (SOC), its computational efficiency is markedly superior, while retaining a good degree of accuracy.
@@
We study the carrier transport properties of ferromagnetic (FM) body-centered cubic (BCC) Fe and face-centered cubic (FCC) Ni when electron phonon interactions are accounted. Our findings indicate that the omission of magnetism in Fe leads to the emergence of imaginary (soft) phonon modes, confirming previous works. As sometimes done in the literature [82, 100, 121, 48], excluding imaginary phonon modes overestimated resistivity due to the large contribution of modes on the onset of being imaginary. In contrast, Ni is a weaker ferromagnet with similar ferromagnetic and non-magnetic phonon band dispersions, but with a very different change of resistivity with temperature. Overall, the inclusion of magnetism reduces resistivity by a factor of three and seven for Fe and Ni, respectively. Finally, we compute the phonon-driven superconducting transition temperature ($T_{\rm c}$) in Fe and Ni and find vanishingly small values, which shows that neither material would be superconducting, even if their Cooper pairs could survive in the FM magnetization field.
This work opens the gate towards accounting for the role of phonons in magnetic systems and possible superconducting states in magnetic materials (such as iron arsenides and possibly cuprates), which are highly relevant for new energy-efficient technologies and energy loss in conducting materials. This work is organized as follows: we first present the theory in Sect. II, details about our computational framework in Sect. III, the interpolation of electron, phonon and electron-phonon quantities in Sect. IV, the calculation of electron-phonon properties in Sect. V, and the final remarks and conclusions in Sect. VI.
@@ II Theory
where $\sigma$ and $\sigma^{\prime}$ are the spinor indices. From the definition of $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$, we can identify two factors. The first is proportional to $\omega_{\mathbf{q}\nu}^{-\frac{1}{2}}$. The second is the spin-dependent deformation potential $D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ which can be computed as [94]
@@
with $m_{\mathrm{u}}$ being the atomic mass constant. $D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ gives a measure of how much a phonon perturbation has changed the internal self-consistent potential felt by the electrons. When taking the absolute value of $|g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})|$, we take the root mean square average over any degenerate manifolds in $m$, $n$ and $\nu$. Note that the definitions of $D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ and $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ are also valid for the case of nonmagnetic compounds with and without SOC. The main property exploited in the nonmagnetic case are the following extra symmetries in the spin indices $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})]^{*}=(-1)^{\sigma+\sigma^{\prime}}g^{\bar{\sigma}^{\prime}\bar{\sigma}}_{mn\nu}(\mathbf{-k},\mathbf{-q})$, with $\bar{\sigma}$ being the opposite spin to $\sigma$. This means that if $\sigma=\sigma^{\prime}$ then $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})]^{*}=g^{\bar{\sigma}^{\prime}\bar{\sigma}}_{mn\nu}(\mathbf{-k},\mathbf{-q})$, and if $\sigma\neq\sigma^{\prime}$ then $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})]^{*}=-g^{\bar{\sigma}^{\prime}\bar{\sigma}}_{mn\nu}(\mathbf{-k},\mathbf{-q})$ (e.g. for heavy elements with spin orbit coupling but no magnetization). These two results are a consequence of time-reversal symmetry, which is absent in magnetic compounds. In practice, performing magnetic calculations with SOC, commonly known as non-collinear magnetic calculations, is computationally expensive (typically by a factor of $2^{3}=8$ for cubic scaling DFT) in systems with several atoms. The collinear approximation consists in separating the two spin channels into up ($\uparrow$) and down ($\downarrow$) without explicit interaction between them, thereby only doubling the cost of the calculations, regardless of the system size. This allows us to define a spin-resolved EPME:
@@
In systems with many valence electrons, we are commonly interested in the deformation potential of several bands. For each spin channel, we can consider
@@
[…]
where $N(\varepsilon_{\textrm{F}})$ is the total density of states at the Fermi energy $\varepsilon_{\rm F}$ and $w_{\mathbf{k}}$ are the $\mathbf{k}$-point integration weights normalized to one. In the non-magnetic case, the sum over $\sigma$ is removed and $\sum_{\mathbf{k}}w_{\mathbf{k}}=2$ due to time-reversal symmetry. We note that another definition exists for the spin-resolved electron-phonon coupling using the partial density of states at the Fermi level $N^{\sigma}(\varepsilon_{\rm F})$ [113, 5]. However, this convention will suffer from divergences in half- and semi-metals, where only one spin channel contributes to the total density of states. The present version should be used systematically for spin polarized cases, taking into account the factor of 2 between $N^{\sigma}$ and $N$ in the unpolarized case.
Another important quantity is the mode- and momentum-resolved phonon linewidth $\gamma_{\mathbf{q}\nu}$
@@
where $w_{\mathbf{q}}$ are the $\mathbf{q}$-point integration weights normalized to one. The Eliashberg spectral function can be used to determine the contribution of specific phonon modes to the total electron-phonon coupling strength defined as
@@
This quantity provides a measure of the overall mass enhancement in a metallic system.
@@
where $w_{\mathbf{q}}$ follow the same integration rules as $w_{\mathbf{k}}$, $f_{m\mathbf{k+q}}^{\sigma}$ is the Fermi-Dirac occupation for spin $\sigma$ and $n_{\mathbf{q\nu}}$ is the Bose-Einstein distribution function. As for previous quantities, the average scattering rate is additive in the collinear approximation to magnetism:
@@
Previous studies have used the lowest-order variational approximation (LOVA) to study the resistivity of magnetic systems [73], which is based on computing the average scattering rate. This is a strong approximation in anisotropic systems or systems with different Fermi sheets [92]. Here, we solve the linearized Boltzmann transport equation (BTE) to calculate the conductivity [95, 27]
@@
which is the usual law of addition for resistivities in parallel, and in this case $\rho_{\alpha\beta}^{\sigma}$ is the resistivity experienced by the electrons in each spin channel. Interestingly, when using the Ziman or LOVA formula, the resistivity becomes additive instead $\rho_{\alpha\beta}^{\uparrow}+\rho_{\alpha\beta}^{\downarrow}$, questioning the validity of Eq. (8) of Ref. [73], see Supplementary Material S1 for more information. When comparing to polycrystalline samples, we use
@@
$\partial_{E_{\beta}}f_{n\mathbf{k}}^{\sigma}=ev_{n\mathbf{k}\beta}^{\sigma}\frac{\partial f^{\sigma}_{n\mathbf{k}}}{\partial\varepsilon^{\sigma}_{n\mathbf{k}}}\tau_{n\mathbf{k}}^{\sigma}+\frac{2\pi\tau_{n\mathbf{k}}^{\sigma}}{\hbar}\sum_{\mathbf{q}m\nu}w_{\mathbf{q}}|g_{mn\nu}^{\sigma}(\mathbf{k},\mathbf{q})|^{2}\\ \times[(n_{\mathbf{q\nu}}+1-f_{n\mathbf{k}}^{\sigma})\delta(\varepsilon_{n\mathbf{k}}^{\sigma}-\varepsilon_{m\mathbf{k+q}}^{\sigma}-\hbar\omega_{\mathbf{q}\nu})\\ \!+\!(n_{\mathbf{q\nu}}\!+\!f_{n\mathbf{k}}^{\sigma})\delta(\varepsilon_{n\mathbf{k}}^{\sigma}\!-\!\varepsilon_{m\mathbf{k+q}}^{\sigma}\!+\!\hbar\omega_{\mathbf{q}\nu})]\partial_{E_{\beta}}f_{m\mathbf{k+q}}^{\sigma}.$ (16)